# Recent developments and research needs in turbulence modeling of hypersonic flows

### v1

@@ I Introduction
The development of hypersonic vehicles, including aircraft, missiles, glide vehicles, reusable launch vehicles, and spacecraft, is at the forefront of aerospace and defense research. Hypersonic vehicles are designed to operate at Mach number greater than $5$ through the planetary atmosphere and they have the potential to transform military capabilities, space exploration, and commercial aviation. A hypersonic vehicle experiences a wide spectrum of flow regimes along its flight trajectory, ranging from subsonic to hypersonic speeds, continuum to free-molecular conditions, and laminar to turbulent flow behavior. These flow conditions are generally characterized using Mach (Ma: vehicle speed/sound speed), Knudsen (Kn: molecular mean free path/reference length), and Reynolds (Re: inertial/viscous forces) numbers. Moreover, high-temperature effects at hypersonic speeds lead to thermochemical non-equilibrium and result in widely varying Damköhler numbers (Da: flow time/chemical reaction times). This review focuses on the modeling of hypersonic turbulent flows in the continuum regime as the vehicle design is generally based on such flow conditions because of the associated extreme aerothermal loads. The complex nature of these flows presents a formidable challenge for accurate physical modeling and numerical analysis and demands sophisticated techniques and significant computational resources.
@@
The complexity of modeling hypersonic turbulent flows relevant to practical applications arises from several unique physical phenomena that are not present or are significantly less pronounced at lower speeds. These include strong compressibility effects; the presence of strong shock waves that lie close to the vehicle body and the associated shock/turbulent boundary layer interactions i.e. SBLIs; significantly high temperatures due to extreme viscous dissipation within the turbulent boundary layers and strong shock waves leading to chemical and thermal non-equilibrium and the associated turbulence/chemistry interactions i.e. TCI; and surface reactions including ablation and the associated ablative particle-flow interaction. Intense aerothermodynamic heating, unsteady pressure loads, and high skin-friction drag are the important global manifestations of the aforementioned complex physical phenomena crucial for engineering design considerations. Additionally, a better understanding of the flow-field in terms of locations of shock waves, expansion fans, turbulent boundary and free-shear layers, and separated regions including shock-induced recirculation bubbles, especially at off-design conditions, is critical for the successful and safe operation of hypersonic vehicles. For example, the shock wave from the engine cowl impinging on the pylon holding a dummy scramjet engine and shock waves impinging on the vertical tail caused severe heating and structural damage to the X-15 aircraft during its final hypersonic flight Anderson (1989). Accurate prediction of these complex hypersonic flow-fields and the engineering quantities of interest (QoIs) is a significantly complex and challenging task for the Computational Fluid Dynamics (CFD) community.
@@ II Governing equations
All terms appearing in the Eqns. (1)-(3) are closed except for those resulting from turbulent fluctuations, represented by the blue underlined terms. These terms describe the average effect of turbulence on the mean flow. The blue underlined terms in Eqns. (2) and (3) include the unclosed terms in the form of the Reynolds stresses $\tau_{ij}=-\overline{\rho u^{{}^{\prime\prime}}_{j}u^{{}^{\prime\prime}}_{i}}$ and turbulent heat flux $q_{T_{j}}=\overline{\rho u^{{}^{\prime\prime}}_{j}h^{{}^{\prime\prime}}}$. The unclosed terms $(\overline{\sigma_{ji}u^{{}^{\prime\prime}}_{i}})$ and $(\overline{\rho u^{{}^{\prime\prime}}_{j}\frac{1}{2}u^{{}^{\prime\prime}}_{i}u^{{}^{\prime\prime}}_{i}})$ in Favre-averaged energy Eqn. (3) characterize the molecular diffusion and turbulent transport of TKE, respectively and they also appear in the Favre-averaged TKE transport equation (see Eqn. (21)). Notably, the gradients of Reynolds stresses in the Favre-averaged momentum Eqn. (2) play a significant role in determining the separation characteristics like separation location and length in case of flows involving shock-induced separations. On the other hand, the gradient of the turbulent heat-flux vector in the Favre-averaged total energy Eqn. (3) is crucial for the correct prediction of wall heat transfer rates. Hence, accurate representation of these turbulent quantities is key to correctly predicting the hypersonic flow field and the engineering QoIs. The goal of turbulence modeling is to develop appropriate closures for these unclosed terms by relating them to known mean flow quantities such as velocity and temperature.
@@ IV.4 Length scale correction
Huang and Coakley Huang and Coakley (1993a) proposed a correction to remedy wall heat transfer overpredictions in the reattachment or shock impingement zone of SBLIs. The correction involves the use of an algebraic length scale which limits the length scale predicted by the two-equation turbulence models, which otherwise would become very large in these regions. For $k$-$\epsilon$ and $k$-$\omega$ models, the correction is given by
@@
Zhang et al. Zhang, Craft, and Iacovides (2022) recently proposed a new turbulent length scale correction for the $\epsilon$-equation taking inspiration from the length-scale modification of Huang and Coakley Huang and Coakley (1993b). The correction is intended to control the growth of turbulence in the SBLI region. This involves the addition of an extra source term to the $\epsilon$-equation,
@@ IV.5 Rapid compression correction
Huang and Coakley Huang and Coakley (1993a) proposed a modification for $k$-$\epsilon$ and $k$-$\omega$ models to improve predictions for flow separation in SBLIs. This modification was made to increase the size of computed separation-bubble regions by ensuring that the turbulent length scale does not change too quickly when undergoing rapid compression. The basic principle of this correction is that the product of the density and turbulent length scale i.e. $\overline{\rho}l$ should remain constant in a uniaxial compression, such as a shock wave, where $l=k^{3/2}/\epsilon=k^{1/2}/\omega$. Applying the continuity equation, the length scale equation can be written as
$\dfrac{1}{l}\dfrac{dl}{dt}=\widetilde{u}_{k,k}.$ (33)
@@ V.1 Flat plate boundary layers
Many one- and two-equation models struggle to accurately predict wall quantities of interest (wall shear stress and wall heat flux) for ZPG flat plate TBLs at hypersonic speeds Aiken et al. (2022); Rumsey (2010); Danis and Durbin (2022); Barone, Parish, and Jordan (2024); particularly for cold wall boundary layers at high Mach and Reynolds numbers Rumsey (2010). In general, RANS models tend to over-predict both wall shear stress and wall heat flux at high Reynolds numbers. Figure 2 illustrates this challenge by showing predictions of two standard RANS models — the Menter SST model and the Spalart–Allmaras model — on a Mach $11$ ZPG hypersonic TBL at a wall temperature-to-recovery ratio of $T_{w}/T_{r}=0.20$, for which comparative DNS data are available from Ref. Huang, Duan, and Choudhari, 2022 and experimental data from the CUBRC facility Gnoffo, Berry, and Van Norman (2011, 2013). For both turbulence models, wall shear stress and heat flux predictions show discrepancies compared to the experiment and DNS that increase as the boundary layer spatially evolves.
@@
Various compressibility corrections have been assessed (see Refs. Rumsey, 2010; Aiken et al., 2022) and proposed (see Refs. Danis and Durbin, 2022; Barone, Parish, and Jordan, 2024) to address these issues. Aiken et al. Aiken et al. (2022) compared the predictions of the SA, SST, and BSL models to DNS data Zhang, Duan, and Choudhari (2018) for ZPG, cold-wall boundary layers for Mach numbers ranging from 2 to 8 with wall-to-recovery-temperature ratios of $0.18$ to 1.0 and concluded that no model accurately predicts QoIs at the wall across this range of conditions. They further concluded that the Zeman compressibility correction tends to improve accuracy for wall QoIs, but does so at the expense of worse wall-normal temperature profiles. Rumsey Rumsey (2010) studies the performance of the $k-\omega$ model over a range of Mach numbers and wall-to-recovery-temperature ratios, and similarly shows degraded performance at high Mach number, cold-wall conditions. Rumsey notes an improved performance of algebraic models like Baldwin-Lomax over standard one- and two-equation models, and further shows that dilatation-dissipation corrections tend to bring $k$-$\omega$ predictions of wall QoIs more inline with the Baldwin-Lomax model. Danis and Durbin propose a compressibility correction for the $k-\omega$ model — which modifies the turbulent production in the $\omega$ equation — to improve wall QoI predictions in hypersonic boundary layers Danis and Durbin (2022). The correction is local, is a function of a friction Mach number and a heat transfer parameter, and shows improved results over a range of boundary layers. Barone et al. Barone, Parish, and Jordan (2024) propose several corrections to the SA model to improve performance in ZPG TBLs including a modified near-wall viscous damping function (informed by hypersonic boundary layer DNS datasets) and a normal Reynolds stress correction. Of note, Ref. Barone, Parish, and Jordan, 2024 highlights that the wall-normal Reynolds stress has an appreciable impact on the wall-normal momentum balance for high-speed boundary layers; this aspect is neglected in traditional RANS models. Specifically, under the assumption that the wall-normal Reynolds stress scales with the freestream dynamic pressure, it is straightforward to show that the ratio of the wall-normal Reynolds stress to pressure scales with the Mach number squared, i.e.,
@@ V.3 Shock/boundary layer interaction and wall heat-transfer
With respect to the joint prediction of wall heat flux and wall shear stress, standard two-equation models tend to massively over-predict wall heat flux near the point of reattachment. This over-prediction is typically worse for $k$-$\epsilon$-based models than for $k$-$\omega$-based models Coakley et al. (1994); Wilcox et al. (1998). Similar to ZPG TBLs, the over-prediction in peak heat flux appears to worsen as the Reynolds number grows Parish et al. (2024a). Length-scale corrections (Sec. IV.4) and variable turbulent Prandtl number models (Sec. IV.8) have shown promise for addressing this issue. Huang et al. Huang and Coakley (1993b) show that length-scale corrections result in a dramatic improvement in wall heat flux predictions for a range of SBLI flows. It is worth highlighting, however, that length-scale corrections additionally reduce wall shear stress; this impact was not assessed in Ref. Huang and Coakley, 1993b. The variable turbulent Prandtl number of Roy and Sinha (see Sec.IV.8) similarly shows promise for addressing over-prediction in wall heat flux. Ref. Roy and Sinha, 2019b shows that an extension of the model for hypersonic flows results in improved predictions for wall heat flux on a variety of hypersonic compression ramps and impinging SBLI cases. Unlike length-scale limiting, the variable turbulent Prandtl number model decouples the turbulent thermal diffusivity from the eddy viscosity. This allows for a reduction in wall heat flux with a minimal impact on the wall shear stress, and vice versa. This approach is supported by experimental data, which suggests that the rate at which wall shear stress rises through an SBLI decouples from the rate at which wall heat flux rises as shock strength increases Schülein (2006). Lastly, we remark that the SA model does not result in the same type of over-predictions in wall heat flux in SBLIs as two-equation models, as seen above. However, the improved performance of the SA model for wall heat flux is at the expense of an under-prediction in wall shear stress Parish et al. (2024a).

### v2

@@ I Introduction
The development of hypersonic vehicles, including aircraft, missiles, glide vehicles, reusable launch vehicles, and spacecraft, is at the forefront of aerospace and defense research. Hypersonic vehicles are designed to operate at Mach number greater than $5$ through the planetary atmosphere and they have the potential to transform military capabilities, space exploration, and commercial aviation. A hypersonic vehicle experiences a wide spectrum of flow regimes along its flight trajectory, ranging from subsonic to hypersonic speeds, continuum to free-molecular conditions, and laminar to turbulent flow behavior. These flow conditions are generally characterized using Mach (M: vehicle speed/sound speed), Knudsen (Kn: molecular mean free path/reference length), and Reynolds (Re: inertial/viscous forces) numbers. Moreover, high-temperature effects at hypersonic speeds lead to thermochemical non-equilibrium and result in widely varying Damköhler numbers (Da: flow time/chemical reaction times). This review focuses on the modeling of hypersonic turbulent flows in the continuum regime as the vehicle design is generally based on such flow conditions because of the associated extreme aerothermal loads. The complex nature of these flows presents a formidable challenge for accurate physical modeling and numerical analysis and demands sophisticated techniques and significant computational resources.
@@
The complexity of modeling hypersonic turbulent flows relevant to practical applications arises from several unique physical phenomena that are not present or are significantly less pronounced at lower speeds. These include strong compressibility effects; the presence of strong shock waves that lie close to the vehicle body and the associated shock/turbulent boundary layer interactions i.e. SBLIs; significantly high temperatures due to extreme viscous dissipation within the turbulent boundary layers and strong shock waves leading to chemical and thermal non-equilibrium and the associated turbulence/chemistry interactions i.e. TCI; and surface reactions including ablation and the associated ablative particle-flow interaction (see Fig. 1). Intense aerothermodynamic heating, unsteady pressure loads, and high skin-friction drag are the important global manifestations of the aforementioned complex physical phenomena crucial for engineering design considerations. Additionally, a better understanding of the flow-field in terms of locations of shock waves, expansion fans, turbulent boundary and free-shear layers, and separated regions including shock-induced recirculation bubbles, especially at off-design conditions, is critical for the successful and safe operation of hypersonic vehicles. For example, the shock wave from the engine cowl impinging on the pylon holding a dummy scramjet engine and shock waves impinging on the vertical tail caused severe heating and structural damage to the X-15 aircraft during its final hypersonic flight Anderson (1989). Accurate prediction of these complex hypersonic flow-fields and the engineering quantities of interest (QoIs) is a significantly complex and challenging task for the Computational Fluid Dynamics (CFD) community.
Figure 1: Complex flow physics around a notional hypersonic flight system. Figure adapted from Ref. Urzay, 2020 with permission from Javier Urzay.
@@ II Governing equations
All terms appearing in the Eqns. (1)-(3) are closed except for those resulting from turbulent fluctuations, represented by the blue underlined terms. These terms describe the average effect of turbulence on the mean flow. The blue underlined terms in Eqns. (2) and (3) include the unclosed terms in the form of the Reynolds stresses $\tau_{ij}=-\overline{\rho u^{{}^{\prime\prime}}_{j}u^{{}^{\prime\prime}}_{i}}$ and turbulent heat flux $q_{T_{j}}=\overline{\rho u^{{}^{\prime\prime}}_{j}h^{{}^{\prime\prime}}}$. The unclosed terms $(D=\overline{\sigma_{ji}u^{{}^{\prime\prime}}_{i}})$ and $(T=\overline{\rho u^{{}^{\prime\prime}}_{j}\frac{1}{2}u^{{}^{\prime\prime}}_{i}u^{{}^{\prime\prime}}_{i}})$ in Favre-averaged energy Eqn. (3) characterize the molecular diffusion and turbulent transport of TKE, respectively and they also appear in the Favre-averaged TKE transport equation (see Eqn. (21)). Notably, the gradients of Reynolds stresses in the Favre-averaged momentum Eqn. (2) play a significant role in determining the separation characteristics like separation location and length in case of flows involving shock-induced separations. On the other hand, the gradient of the turbulent heat-flux vector in the Favre-averaged total energy Eqn. (3) is crucial for the correct prediction of wall heat transfer rates. Hence, accurate representation of these turbulent quantities is key to correctly predicting the hypersonic flow field and the engineering QoIs. The goal of turbulence modeling is to develop appropriate closures for these unclosed terms by relating them to known mean flow quantities such as velocity and temperature.
Models for the Reynolds stresses $\tau_{ij}$ fall into two classes: “eddy viscosity” relations between the stresses and the mean velocity gradients at the same point in space; and models of the further unknowns in the exact partial differential transport equations for the individual Reynolds stress components. Eddy viscosity models are based on an analogy to laminar viscous transport given by Eqn. (7) introducing eddy viscosity $\mu_{T}$ as an unknown and the exact formulation of $\mu_{T}$ is model dependent. The turbulent heat flux vector, on the other hand, is generally modeled in analogy to the laminar heat flux vector given by Eqn. (8) using a gradient diffusion hypothesis with a turbulent Prandtl number $Pr_{t}$ as an unknown model coefficient. The turbulent Prandtl number $Pr_{t}$ is typically set to a constant value, however, models that vary $Pr_{t}$ in a flow field using algebraic relations or partial differential equations also exist. The unknown molecular diffusion $D$ and turbulent transport $T$ (also known as triple velocity correlation) terms are also modeled using a gradient diffusion hypothesis albeit for the TKE which introduces unknown model coefficients the values of which depend on the specific turbulence model.
@@ IV.4 Length scale correction
Vuong and Coakley Vuong and Coakley (1987) and Huang and Coakley Huang and Coakley (1993a) proposed a correction to remedy wall heat transfer overpredictions in the reattachment or shock impingement zone of SBLIs. The correction involves the use of an algebraic length scale which limits the length scale predicted by the two-equation turbulence models, which otherwise would become very large in these regions. For $k$-$\epsilon$ and $k$-$\omega$ models, the correction is given by
@@
Zhang et al. Zhang, Craft, and Iacovides (2022) recently proposed a new turbulent length scale correction for the $\epsilon$-equation taking inspiration from the length-scale modification discussed above. The correction is intended to control the growth of turbulence in the SBLI region. This involves the addition of an extra source term to the $\epsilon$-equation,
@@ IV.5 Rapid compression correction
Vuong and Coakley Vuong and Coakley (1987) and Huang and Coakley Huang and Coakley (1993a) proposed a modification for $k$-$\epsilon$ and $k$-$\omega$ models to improve predictions for flow separation in SBLIs. This modification was made to increase the size of computed separation-bubble regions by ensuring that the turbulent length scale does not change too quickly when undergoing rapid compression. The basic principle of this correction is that the product of the density and turbulent length scale $\overline{\rho}l$, where $l=k^{3/2}/\epsilon=k^{1/2}/\omega$, should remain constant in a uniaxial compression, such as a shock wave, i.e.
$\dfrac{d\overline{\rho}l}{dt}=0.$ (33)
Applying the continuity equation, $(1/\overline{\rho})d\overline{\rho}/dt=-\partial\widetilde{u}_{k}/\partial x_{k}$, to the above equation, the length scale equation can be written as
$\dfrac{1}{l}\dfrac{dl}{dt}=\dfrac{\partial\widetilde{u}_{k}}{\partial x_{k}}.$ (34)
@@ IV.7 Diffusion term corrections
IV.8 $\omega$-equation correction
Danis and Durbin Danis and Durbin (2022) proposed a compressibility correction for two-equation $k$-$\omega$ models which explicitly takes the effects of high Mach numbers and wall cooling in account. The compressibility correction is heuristic and aims to improve the predictions of the $k$-$\omega$ models that inaccurately predict high values of eddy viscosity and low mean velocities, as wall cooling increases. The correct profiles of eddy viscosity are achieved by modifying the closure coefficients of the production and destruction terms in the $\omega$-equation while the $k$-equation remains unchanged.
The baseline production and destruction terms in the $\omega$-equation are multiplied with a compressibility correction function $f$ given by
$f=(1-\phi)\exp{-c_{1}M^{*2}_{\tau}-c_{2}B^{*}_{q}})+\phi\exp{c_{3}M^{*}_{\tau}+c_{4}B^{*}_{q}},$ (39)
where
$\begin{split}c_{1}=60,c_{2}=5,c_{3}=\dfrac{15}{4}\exp{\left(-\dfrac{M}{10}\right)},\\ c_{4}=-\dfrac{21}{4}\exp{\left(\dfrac{M}{5}\right)},\phi=\tanh{\left(\dfrac{3}{4}M\right)}.\end{split}$ (40)
Here, $M$ is the local Mach number, friction Mach number $M^{*}_{\tau}=\widetilde{u}^{*}/a$ with $\widetilde{u}^{*}$ being the function of the total local stress $\tau$ given by $\widetilde{u}^{*}=\sqrt{\tau/\overline{\rho}}$ and the heat flux parameter $B^{*}_{q}=q/(\overline{\rho}C_{p}\widetilde{T}\widetilde{u}^{*})$ with $q$ being the magnitude of the local heat flux components $q_{x_{j}}$ that are constructed from the boundary-layer approximations to the enthalpy equation and represent the effects of the molecular and turbulent diffusion of the mean enthalpy as well as the dissipative heating.
The form of $f$ is based on extensive numerical experiments. The first term in Eqn. (39) is active below the log layer and it controls the $y$ intercept of the log law while the second term is responsible for the log-law slope. The correction function $f$ is deactivated i.e $f\rightarrow 1$ at low speed as $M_{\infty}\rightarrow 0$ and $Tw/T_{r}\rightarrow 1$. The correction is, however, not Galilean invariant as it depends on the local Mach number $M$.
The correction is validated using the DNS database Zhang, Duan, and Choudhari (2018); Huang et al. (2020) for ZPG cases for a range of Mach numbers between $2$ and $14$ and wall-to-recovery temperature ratios ($T_{w}/T_{r}$) between $0.18$ and $1$ as well as Mach $5$ favorable and adverse pressure gradient cases Nicholson et al. (2021a, b) with $T_{w}/T_{r}=0.91$. The correction collapses the eddy viscosity profiles on the incompressible data giving significantly improved velocity, temperature, skin friction, and wall heat transfer coefficient profiles than the baseline $k$-$\omega$ models.
@@ V.1 Flat plate boundary layers
Many one- and two-equation models struggle to accurately predict wall quantities of interest (wall shear stress and wall heat flux) for ZPG flat plate TBLs at hypersonic speeds Aiken et al. (2022); Rumsey (2010); Danis and Durbin (2022); Barone, Parish, and Jordan (2024a); particularly for cold wall boundary layers at high Mach and Reynolds numbers Rumsey (2010). In general, RANS models tend to over-predict both wall shear stress and wall heat flux at high Reynolds numbers. Figure 3 illustrates this challenge by showing predictions of two standard RANS models — the Menter SST model and the Spalart–Allmaras model — on a Mach $11$ ZPG hypersonic TBL at a wall temperature-to-recovery ratio of $T_{w}/T_{r}=0.20$, for which comparative DNS data are available from Ref. Huang, Duan, and Choudhari, 2022 and experimental data from the CUBRC facility Gnoffo, Berry, and Van Norman (2011, 2013). For both turbulence models, wall shear stress and heat flux predictions show discrepancies compared to the experiment and DNS that increase as the boundary layer spatially evolves.
@@
Various compressibility corrections have been assessed (see Refs. Rumsey, 2010; Aiken et al., 2022) and proposed (see Refs. Danis and Durbin, 2022; Barone, Parish, and Jordan, 2024a) to address these issues. Aiken et al. Aiken et al. (2022) compared the predictions of the SA, SST, and BSL models to DNS data Zhang, Duan, and Choudhari (2018) for ZPG, cold-wall boundary layers for Mach numbers ranging from 2 to 8 with wall-to-recovery-temperature ratios of $0.18$ to 1.0 and concluded that no model accurately predicts QoIs at the wall across this range of conditions. They further concluded that the Zeman compressibility correction tends to improve accuracy for wall QoIs, but does so at the expense of worse wall-normal temperature profiles. Rumsey Rumsey (2010) studies the performance of the $k-\omega$ model over a range of Mach numbers and wall-to-recovery-temperature ratios, and similarly shows degraded performance at high Mach number, cold-wall conditions. Rumsey notes an improved performance of algebraic models like Baldwin-Lomax over standard one- and two-equation models, and further shows that dilatation-dissipation corrections tend to bring $k$-$\omega$ predictions of wall QoIs more inline with the Baldwin-Lomax model. Barone et al. Barone, Parish, and Jordan (2024a) propose several corrections to the SA model to improve performance in ZPG TBLs including a modified near-wall viscous damping function (informed by hypersonic boundary layer DNS datasets) and a normal Reynolds stress correction. Of note, Ref. Barone, Parish, and Jordan, 2024a highlights that the wall-normal Reynolds stress has an appreciable impact on the wall-normal momentum balance for high-speed boundary layers; this aspect is neglected in traditional RANS models. Specifically, under the assumption that the wall-normal Reynolds stress scales with the freestream dynamic pressure, it is straightforward to show that the ratio of the wall-normal Reynolds stress to pressure scales with the Mach number squared, i.e.,
@@ V.3 Shock/boundary layer interaction and wall heat-transfer
With respect to the joint prediction of wall heat flux and wall shear stress, standard two-equation models tend to massively over-predict wall heat flux near the point of reattachment. This over-prediction is typically worse for $k$-$\epsilon$-based models than for $k$-$\omega$-based models Coakley et al. (1994); Wilcox et al. (1998). Similar to ZPG TBLs, the over-prediction in peak heat flux appears to worsen as the Reynolds number grows Parish et al. (2024a). Length-scale corrections (Sec. IV.4) and variable turbulent Prandtl number models (Sec. IV.9) have shown promise for addressing this issue. Huang et al. Huang and Coakley (1993b) show that length-scale corrections result in a dramatic improvement in wall heat flux predictions for a range of SBLI flows. It is worth highlighting, however, that length-scale corrections additionally reduce wall shear stress; this impact was not assessed in Ref. Huang and Coakley, 1993b. The variable turbulent Prandtl number of Roy and Sinha (see Sec.IV.9) similarly shows promise for addressing over-prediction in wall heat flux. Ref. Roy and Sinha, 2019b shows that an extension of the model for hypersonic flows results in improved predictions for wall heat flux on a variety of hypersonic compression ramps and impinging SBLI cases. Unlike length-scale limiting, the variable turbulent Prandtl number model decouples the turbulent thermal diffusivity from the eddy viscosity. This allows for a reduction in wall heat flux with a minimal impact on the wall shear stress, and vice versa. This approach is supported by experimental data, which suggests that the rate at which wall shear stress rises through an SBLI decouples from the rate at which wall heat flux rises as shock strength increases Schülein (2006). Lastly, we remark that the SA model does not result in the same type of over-predictions in wall heat flux in SBLIs as two-equation models, as seen above. However, the improved performance of the SA model for wall heat flux is at the expense of an under-prediction in wall shear stress Parish et al. (2024a).